OpenStudy (anonymous):

What is the maximum number of relative extrema contained in the graph of this function: f(x)=3x^4-x^2+4x-2

OpenStudy (anonymous):

hint: take the derivative of the function. then examine how many zeros that function can have, namely the max number of real zeros. that will be your answer.

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

what's the dreivative?

OpenStudy (anonymous):

derivative

OpenStudy (anonymous):

f(x)=12x^3-2x+4

OpenStudy (anonymous):

all right. what degree polynomial is the derivative?

OpenStudy (anonymous):

cubic

OpenStudy (anonymous):

how many zeros will it have?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

really? have you heard of the fundamental theorem of algebra?

OpenStudy (anonymous):

Yeah? it has at least one complex root

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html have a look at this...

OpenStudy (anonymous):

so 3

OpenStudy (anonymous):

@pgpilot326 where'd you go!

OpenStudy (anonymous):

right here... great. does that make sense?

OpenStudy (anonymous):

Yeah,so how do I find max number of relative extrema

OpenStudy (anonymous):

each relative extrema will occur at a zero of the derivative function. so if your function is a polynomial then the max number will be 1 less than it's degree.

OpenStudy (anonymous):

So the max number is 3?

OpenStudy (anonymous):

yep and the minimum number is 1. that's because when a polynomial has real coefficients, complex roots occur in conjugate pairs. since 3 is the max, 3-2 = 1 would be the min. (number of relative extrema won't be below 0).

OpenStudy (anonymous):

Ok that makes a lot of sense! Would you mind helping on one more? It is pretty easy identifying a polynomial from a multiple choice list. Just be check my answer

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

WHich is a polynomial: f(x)=4/x - 11x^2 f(x)= -2.1^7+3x^2+17/2 x ---- my answer f(x)= -1/5x^2 +4.5x + 2^x f(x)= -3x^4+8x^2-2sqrtx + 1

OpenStudy (anonymous):

a polynomial must have non-negative, integer exponents.

OpenStudy (anonymous):

Ok my answer is correct then

OpenStudy (anonymous):

so in your answer, is the last term \[\frac{ 17 }{ 2x }\text{ or }\frac{ 17 }{ 2 }x\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you're welcome!